OpenStudy (anonymous):

Please solve 4/7=x+3/21

OpenStudy (anonymous):

x = 159/4 = 39.75.

OpenStudy (anonymous):

Cross multiply.

OpenStudy (anonymous):

A. x = 91 B. 33 3/4 C. x = 15 D. x = 9

OpenStudy (anonymous):

Tell me what answer you get. @kk54

OpenStudy (anonymous):

I got 3/7 but that's not one of the answer choices

OpenStudy (anonymous):

I did exactly that.

OpenStudy (tester97):

\[\frac{ 4 }{ 7 }=x+\frac{ 3 }{ 21 }\] \[\frac{ 3 }{ 21 } simplified~\to~\frac{ 1 }{ 7 }\] \[\frac{ 4 }{ 7 }=x+\frac{ 1 }{ 7 }\] \[subtract \frac{ 1 }{ 7 }\] \[\frac{ 3 }{ 7 }=x\]

OpenStudy (anonymous):

The answer is D. x=9

OpenStudy (tester97):

Those answer choices dont make sense with that equation.

OpenStudy (anonymous):

@vzfreakz is correct..!